OpenStudy (anonymous):

Simplify the rational expression. Find all numbers that must be excluded from the domain of the simplified rational expression. 3x+2/15x ^{2}+16x+4

OpenStudy (anonymous):

To find this you need to figure out where the denominator would equal zero because we can't divide by zero. I used the quadratic equation to factor the denominator as follows:\[-b \pm \sqrt{b ^{2}}\-2a] a = 15; b = 16; c = 4 There are two real solutions where the denominator equals zero at x=-3/10 and x=-2/3, so these two points would be excluded from the domain.